Additional Problem #3

Price Discrimination Cost Structure MC = 20 Revenue Structure P = 80 – Q

Competitive

Imperfect Comp.

1st Degree Price Discrimination

2nd Degree Price Discrimination $50, $30 & $20

Quantity __________ __________ xxxxxxxxx xxxxxxxxx

Price __________ __________ xxxxxxxxx xxxxxxxxx

Consumer Surplus __________ __________ ________ ________

Producer Surplus __________ __________ ________ ________

Total Surplus __________ __________ ________ ________

Dead-weight loss __________ __________ ________ ________
Answer #1

mc-20 P= 80-9 TR= Sog-g² MR = 80-29. 1st Degree Price Discrimination it is practice To Charge different price for each and evCost and Revenue. 80 amc=20 20 → D = P = AR=MR bot so quantity quantity : 60 Price : 20 Consumer Susius : 0 why Because of fi- 2nd degree Price discrimination: His Practice of Charging different price for different Blocks of onit Purchased. Cost andLost and Revenue. quantity = 50 Price = 30 Consumer Surhus, ₂ (50) (50) P=30 1250 Ik - MG 50 60 so quantity Producer Surfius.